Exploring photoproduced meson states that decay K^{+}K^{−}π^{0} with an emphasis on events where the K^{+}K^{−} pair are from a_{0} decay

A. Gardner

Abstract

Mesons that have K^{+}K^{−}π^{0} decay products are potentially interesting as exotic states, however many non-exotics have the same decay products. On the way towards uncovering exotic meson states, a major goal of the GlueX experiment is to map the non-exotic light mesons into their respective nonets. This endeavor requires that methods utilized can sufficiently measure well known states. As a step towards uncovering the richness possessed for mesons that have K^{+}K^{−}π^{0} decay products is to search for low mass states decaying a_{0}π^{0} that include K^{+}K^{−}π^{0} non-resonant contributions. Utilizing GlueX data for K^{+}K^{−}π^{0} photoproduction from the proton I will show a sharp peak in invariant mass of K^{+}K^{−}π^{0} near the f_{1}(1285) and η(1295) mass along with angular distribution of the decay products.